OpenStudy (anonymous):

What is the surface area of a rectangular prism with length 15 m, width 9 m, and height 6 m? 810 m2 450 m2 558 m2 279 m2 @cp9454

OpenStudy (cp9454):

558

OpenStudy (nurali):

A=2(wl+hl+hw)= A=2·(9·15+6·15+6·9) A=558m²